# A discussion on the symmetry of relativistic Vlasov gas and its accretion in Kerr-Newman black hole

### v1

@@ Abstract
We investigate the kinetic properties of collisionless Vlasov gas in Kerr-Newman spacetime, focusing on how spacetime symmetries constrain the distribution functions. By establishing the general relationship between distribution functions and constants of motion (mass $m$, energy $E$, angular momentum $L_{z}$, Carter constant $L$) along with $L$’s conjugate configuration variable $Q^{3}$, we derive the analytical framework for particle dynamics in stationary axisymmetric spacetimes. Within the Locally Non-Rotating Frame , we compute key observables—particle number density, energy density, principal pressures, and accretion rates—resolving coordinate divergences near the event horizon through geometric rescaling. For a plasma obeying the Jüttner distribution, asymptotic expressions at spatial infinity are explicitly calculated. Numerical analyses reveal: (1) mass and energy accretion rates exhibit nearly identical parametric dependencies on black hole parameters; (2) extremal Kerr-Newman black holes show monotonic enhancement in mass, energy, and angular momentum accretion rates with increasing $a$, contrasting with Kerr black hole behavior. These results provide new insights into accretion mechanisms in extreme spacetime geometries.
@@ I Introduction
Research on accretion can be traced back to the early work of Lyttleton, Hoyle and Bondi, who first provided a relativistic solution for perfect fluids Hoyle(1939) ; Bondi(1944) ; Bondi(1952) , and these were later extended to general relativity by Michel Michel(1972) . When particle collisions become infrequent, the perfect fluid approximation breaks down. In such collisionless regimes, kinetic theory governed by the Vlasov equation provides a physically viable alternative.
@@
The theory of gas dynamics was first introduced to special relativity by Jüttner Juttner(1911) , and was later extended by Synge, Tauber, Israel, and others, and gradually incorporated into general relativity Synge(1934) ; Tauber(1961) ; Israel(1963) ; Ehlers(1971) ; Ehlers(1973) ; Cercignani(2002) . The study of kinetic theory of gases helps to understand the properties of accretion disks and the jet launching mechanism, such as reconstructing images of supermassive black holes at the center of M87 galaxy Narayan(1995) . In addition, observations indicate that the gas surrounding M87 and Sgr A* is nearly collisionless and magnetized Akiyama(2019) ; Akiyama(2022) . For studies on collisionless gas models in galaxies and galaxy clusters, see Refs. Rein(2014) ; Hadvzic(2021) .
@@
Recent years have witnessed extensive investigation of relativistic Vlasov gases. Sarbach and his collaborators established a foundational paradigm, first analyzing relativistic kinetic gases and their distribution functions, and then constructing a rigorous geometric framework on both tangent and cotangent bundles Sarbach(201303) ; Sarbach(201309) ; Sarbach(201311) ; Sarbach(2022) . Rioseco implemented this framework numerically in Schwarzschild spacetime, revealing stark contrasts between Vlasov gas accretion and perfect fluid accretion. These results partially explain the observed accretion rates substantially below Bondi predictions Rioseco-Sarbach(20171) ; Rioseco-Sarbach(20172) . Gamboa Gamboa-Sarbach(2021) and Liao Liao(2022) incorporated angular momentum distributions, addressing phenomena like low luminosity. Mach et al. further modeled accretion onto moving Schwarzschild black holes using Vlasov gases for dark matter simulations, achieving robust results via Monte Carlo methods Mach(20211) ; Mach(20212) ; Mach(2022) ; Mach(2023) ; Mach(2024) .
@@
For non-Schwarzschild black holes, Cieślik Cieslik-Mach(2020) and Li Li(2025) demonstrated that black hole charge suppresses mass accretion rates, and considered both charged black holes and charged gases. In stationary axisymmetric spacetimes, they developed Kerr accretion models for gas confined to the equatorial plane Cieslik-Mach(2022) ; Mach-Odrzywolek(2023) and unbound configurations Li(2023) . Discussions on the phase space structure of gas in stationary axisymmetric spacetimes and related applications can be found in Refs. Rioseco-Sarbach(2023) ; Rein(2014) ; Rein(2023) ; Amesl(2016) ; Cai(2023) .
@@
This paper is organized as follows. Sec. II discusses how the Vlasov equation and spacetime symmetries constrain the distribution function in Kerr-Newman (KN) spacetime. Sec. III analyzes charged particle trajectories. Sec. IV derives physical quantities: particle number density, pressures, accretion rates, and asymptotic behavior. Sec. V presents numerically computed profiles of these physical quantities. Finally, Sec. VI offers a concluding summary of the work.
@@
We adopt these conventions: a 4-dimensional Lorentzian manifold is denoted by $(M,g)$, with local coordinates $x^{\mu}$, Greek letters $\mu,\nu,...\in\{0,1,2,3\}$. The metric signature is (,+,+,+)(-,+,+,+)( - , + , + , + ), with geometric units $G=c=1$. The tangent bundle is $TM$, with $T_{x}M$ being the tangent space at $x\in M$. The cotangent bundle is $T^{*}M$, with $T^{*}_{x}M$ the cotangent space at $x$. Lie derivatives along vector field $v$ are denoted $\mathscr{L}_{v}$.
@@ II.1 One-Particle Distribution Function and Its Invariance
In spacetime $(M,g)$, the dynamical state of a particle with mass $m$ is specified by a position $x\in M$ and physical momentum $p_{a}\in T^{*}_{x}M$. The future mass shell is
$\Gamma^{+}_{m}:=\left\{(x,p_{a})\in T^{*}M:x\in M,p_{a}\in F^{+}{x}(m)\right\},$ (1)
@@
$f:\Gamma^{+}\to\mathbb{R},$ (3)
where $\Gamma^{+}\equiv\displaystyle\cup_{m>0}\Gamma^{+}_{m}$. The particle current density $J_{\mu}$ and energy-momentum tensor $T_{\mu\nu}$ are
$J_{\mu}|_{x}$ $:=\int_{F^{+}_{x}}p_{\mu}f\mathrm{dvol}_{x}(p),$ $T_{\mu\nu}|_{x}$ $:=\int_{F^{+}_{x}}p_{\mu}p_{\nu}f\mathrm{dvol}_{x}(p),$ (4)
where $\mathrm{dvol}_{x}(p)$ represents the volume element on $F^{+}_{x}\equiv\displaystyle\cup_{m>0}F^{+}_{x}(m)$.
For free particles (influenced solely by the spacetime background) with mass $m$ and charge $q$ in Kerr-Newman (KN) spacetime with electromagnetic field $F_{ab}=(\mathrm{d}A)_{ab}$, the equations of motion for a trajectory $C(\tau)$ are
@@
$\hat{C}(\tau):=\left\{(x^{\mu}(\tau),p_{\nu}(\tau)):p_{\mu}(\tau)\equiv mg_{\mu\nu}|_{x(\tau)}T^{\nu}(\tau)\right\},$ (6)
@@
Next we consider Killing vector fields on $M$. Let $\phi:M\to M$ be the isometry generated by a Killing vector field $K$. The original trajectory $C(\tau)$ naturally induces a transformed trajectory $\tilde{C}(\tau)\equiv\phi\circ C(\tau)$, with tangent vector $\tilde{T}^{c}|_{\phi(x)}=\phi_{*}[T^{c}|_{x}]$, which preserves the dynamical equations due to the invariance properties,
@@
Beyond these symmetry considerations, we impose no additional constraints on the distribution function $f:\Gamma^{+}\to\mathbb{R}$.
@@ II.2 Distribution Function in The Phase Space
where $\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\varDelta=r^{2}-2\mathcal{M}r+a^{2}+Q^{2},\varSigma^{2}=(r^{2}+a^{2})^{2}-\varDelta a^{2}\sin^{2}\theta$, and $a=\frac{J}{M}$. The electromagnetic vector potential is given by
@@
Killing vectors $\xi^{a}=\left(\frac{\partial\ }{\partial t}\right)^{a},\psi^{a}=\left(\frac{\partial\ }{\partial\varphi}\right)^{a}$ and Killing tensorsCarter(1973) ; Wald(1984) $g_{ab},K_{ab}\equiv 2\rho^{2}l_{(a}k_{b)}+r^{2}g_{ab}$ (where $l^{a}\equiv\frac{1}{\ \Delta\ }\left[(r^{2}+a^{2})\xi^{a}+a\psi^{a}+\Delta\left(\frac{\partial\ }{\partial r}\right)^{a}\right],k^{a}\equiv\frac{1}{2\rho^{2}}\left[(r^{2}+a^{2})\xi^{a}+a\psi^{a}-\Delta\left(\frac{\partial\ }{\partial r}\right)^{a}\right]$ ) provide constants of motion
$E$ $:=-\pi_{a}\xi^{a},$ $L_{z}$ $:=\pi_{a}\psi^{a},$ $H$ $:=\frac{1}{2}g_{ab}p^{a}p^{b},$ $L^{2}$ $:=K_{ab}p^{a}p^{b},$ (12)
@@
Using the these constants of motion as the canonical momentum coordinates,
@@
the conjugate configuration variables are defined via
$Q^{\mu}:=\frac{\partial S}{\partial P_{\mu}},$ (14)
@@
where $\pi(K)\equiv\left[\pi_{\mu}(\mathrm{d}x^{\mu})_{A}\right]\left[K^{\mu}\left(\frac{\partial\ }{\partial x}^{\mu}\right)^{A}\right]$ and $K^{a}\equiv K^{\mu}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{a}$ is a Killing vector on $M$. This yields the explicit form of the Vlasov equation Eq. (7):
@@ III Particle Trajectory Analysis
The equations of motion Eq. (12) for charged particles in KN spacetime are
@@
where $\kappa\equiv qQ$. To facilitate subsequent discussions, we now define several dimensionless parameters:
@@
The angular motion Eq. (20b) yields the following parametrized system
@@
where $\bar{L}\equiv\sqrt{L^{2}-a^{2}\cos^{2}\theta}$. The radial motion equation (20a) can be reformulated as
@@
so, $E\geq 1$ is necessary.
At a given spatial position $(r,\theta)$, the angular motion Eq. (20b) requires $\bar{L}\geq 0$, while the radial motion Eq. (24) demands $R(r)\geq 0$. These constraints define the parameter ranges $D_{\text{abs}}$ for particles absorbed by the black hole and $D_{\text{scat}}$ for scattered particles:
$D_{\text{abs}}$ $\equiv D_{\sigma}\times D_{E\text{abs}}\times D_{\bar{L}\text{abs}}=[0,2\pi]\times[1,+\infty)\times(0,\bar{L}_{c}),$ (26a)
$D_{\text{scat}}$ $\equiv D_{\sigma}\times D_{E\text{scat}}\times D_{\bar{L}\text{scat}}=[0,2\pi]\times[E_{c},+\infty)\times(\bar{L}_{c},\bar{L}_{\text{max}}),$ (26b)
where the parameter $\bar{L}$ (rather than $L$) is adopted due to its convenience when studying specific spatial positions. The upper bound $\bar{L}_{\mathrm{max}}=\bar{L}_{\mathrm{max}}(r,\theta;\sigma,E)$ is determined by the $R(r,\theta;\sigma,E,\bar{L})=0$,
@@
The critical parameters $E_{c}-r$ and $\bar{L}_{c}-E$ can be determined by solving the periastron equation ($\dot{r}=0$): $R=0$ and $R^{\prime}\equiv\frac{\partial R}{\partial r}=0$. Simplifying under the condition $02rE-\kappa>$ 0 yields:
@@
Decoupling $E$ and $\bar{L}$ (involving roots of quartic equations) gives formal expressions $E=f_{E}(a,Q,r,\theta,\sigma,q)$, and $\bar{L}=f_{\bar{L}}(a,Q,r,\theta,\sigma,q)$. Numerical analysis reveals two critical radii for $r>r_{H}=1+\sqrt{1-(a^{2}+Q^{2})}$: a singularity $r_{\text{ph}}=r_{\text{ph}}(a,Q,\theta,\sigma)$ where $f_{E}\to+\infty$ and a root $r_{\text{mb}}=r_{\text{mb}}(a,Q,\theta,\sigma,q)$ where $f_{E}=1$. The monotonicity of $f_{E}-r$ over $(r_{\text{ph}},r_{\text{mb}})$ permits the formal derivation of $\bar{L}=g_{\bar{L}}(a,Q,\theta,E,\sigma,q)$
$E_{c}$ =$=\begin{cases}+\infty&\text{for }r\leq r_{\text{ph}};\\ f_{E}(a,Q,r,\theta,\sigma,q),&\text{for }r_{\text{ph}}<r<r_{\text{mb}};\\ 1,&\text{for }r\geq r_{\text{mb}}.\end{cases},$ $\bar{L}_{c}$ =$=\begin{cases}g_{\bar{L}}(a,Q,\theta,E,\sigma,q)&\text{for }E>1;\\ \bar{L}_{\text{mb}}&\text{for }E=1.\end{cases}$ (29)
where $\bar{L}_{\text{mb}}\equiv f_{\bar{L}}|_{r=r_{\text{mb}}}$. Numerical examples are shown in Fig. 1, while analytical derivations of $E_{c}$ and $\bar{L}_{c}$ are detailed in Appendix B.
@@
Figure 1: For a KN black hole ($M=1$, $a=0.1$, $Q=0.1$), critical parameters (a) $E_{c}$ at the periastron radii $r$ and (b) $L_{c}-E$ on the equatorial plane ($\theta=\pi/2$) for various $\kappa$, with $\sigma=\pi/2$.
Fig. 1a shows $E_{c}$ at periastron radii. For fixed $\kappa$, $E_{c}(r)$ has a vertical asymptote at $r=r_{\text{ph}}$ and $E_{c}(r_{\text{mb}})=1$. When $r\in(r_{\text{ph}},r_{\text{mb}})$, $E_{c}$ increases as $r$ decreases. No scattering occurs for $r<r_{\text{ph}}$, while any energy $E_{c}\geq 1$ permits scattering when $r>r_{\text{mb}}$. Fig. 1b shows that $L_{c}$ grows approximately linearly with $E$ for fixed $\kappa$, with curves shifting downward as $\kappa$ increases.

### v2

@@ Abstract
We investigate the kinetic properties of collisionless Vlasov gas in the Kerr-Newman spacetime, focusing on how the spacetime symmetries constrain the distribution functions. By establishing the general relationship between the distribution functions and the constants of motion (mass $m$, energy $E$, angular momentum $L_{z}$, Carter constant $L$) along with $L$’s conjugate configuration variable $Q^{3}$, we derive the analytical framework for particle dynamics in stationary axisymmetric spacetimes. Within the Locally Non-Rotating Frame, we compute key observables—the particle number density, the energy density, the principal pressures, and the accretion rates—resolving coordinate divergences near the event horizon through geometric rescaling. For a plasma obeying the Jüttner distribution, asymptotic expressions at spatial infinity are explicitly calculated. Numerical analyses reveal: (1) the mass and energy accretion rates exhibit nearly identical parametric dependencies on the black hole parameters; (2) extremal Kerr-Newman black holes show monotonic enhancement in the mass, energy, and angular momentum accretion rates with increasing $a$, contrasting with the Kerr black hole behavior. These results provide new insights into accretion mechanisms in extreme spacetime geometries.
@@ I Introduction
Research on accretion can be traced back to the early work of Lyttleton, Hoyle and Bondi, who first provided a relativistic solution for perfect fluids Hoyle(1939) ; Bondi(1944) ; Bondi(1952) , and these studies were later extended to general relativity by Michel Michel(1972) . When particle collisions become infrequent, the perfect fluid approximation breaks down. In such collisionless regimes, kinetic theory governed by the Vlasov equation provides a physically viable alternative.
@@
The theory of gas dynamics was first introduced into special relativity by Jüttner Juttner(1911) , and was later extended by Synge, Tauber, Israel, and others, gradually incorporating into general relativity Synge(1934) ; Tauber(1961) ; Israel(1963) ; Ehlers(1971) ; Ehlers(1973) ; Cercignani(2002) . The study of the kinetic theory of gases helps to understand the properties of accretion disks and the jet launching mechanism, such as reconstructing images of the supermassive black holes at the center of the M87 galaxy Narayan(1995) . In addition, observations indicate that the gas surrounding M87 and Sgr A* is nearly collisionless and magnetized Akiyama(2019) ; Akiyama(2022) . For studies on collisionless gas models in galaxies and galaxy clusters, see Refs. Rein(2014) ; Hadvzic(2021) .
@@
Recent years have witnessed extensive investigation of relativistic Vlasov gases. Sarbach and his collaborators established a foundational paradigm, first analyzing relativistic kinetic gases and their distribution functions, and then constructing a rigorous geometric framework on the tangent and cotangent bundles Sarbach(201303) ; Sarbach(201309) ; Sarbach(201311) ; Sarbach(2022) . Rioseco implemented the framework numerically in the Schwarzschild spacetime, revealing stark contrasts between Vlasov gas accretion and perfect fluid accretion. These results partially explain the observed accretion rates substantially below the Bondi predictions Rioseco-Sarbach(20171) ; Rioseco-Sarbach(20172) . Gamboa Gamboa-Sarbach(2021) and Liao Liao(2022) incorporated angular momentum distributions, addressing phenomena such as low luminosity. Mach et al. further modeled accretion onto moving Schwarzschild black holes using Vlasov gases for dark matter simulations, achieving robust results via Monte Carlo methods Mach(20211) ; Mach(20212) ; Mach(2022) ; Mach(2023) ; Mach(2024) .
@@
For non-Schwarzschild black holes, Cieślik Cieslik-Mach(2020) and Li Li(2025) demonstrated that black hole charge suppresses the mass accretion rates, and considered both charged black holes and charged gases. In stationary axisymmetric spacetimes, they developed Kerr accretion models for gas confined to the equatorial plane Cieslik-Mach(2022) ; Mach-Odrzywolek(2023) and unbound configurations Li(2023) . Discussions regarding the phase space structure of gas in stationary axisymmetric spacetimes and related applications can be found in Refs. Rioseco-Sarbach(2023) ; Rein(2014) ; Rein(2023) ; Amesl(2016) ; Cai(2023) .
@@
This paper is organized as follows. Sec. II discusses how the Vlasov equation and spacetime symmetries constrain the distribution function in Kerr-Newman (KN) spacetime. Sec. III analyzes charged particle trajectories. Sec. IV derives physical quantities: the particle number density, pressures, accretion rates, and asymptotic behavior. Sec. V presents numerically computed profiles of these physical quantities. Finally, Sec. VI offers a concluding summary of the work.
@@
We adopt these conventions: A 4-dimensional Lorentzian manifold is denoted by $(M,g)$, with local coordinates $x^{\mu}$ (where Greek letters $\mu,\nu,...\in\{0,1,2,3\}$). The metric signature is (,+,+,+)(-,+,+,+)( - , + , + , + ), with geometric units $G=c=1$. The tangent bundle is $TM$, with $T_{x}M$ denoting the tangent space at $x\in M$. The cotangent bundle is $T^{*}M$, with $T^{*}_{x}M$ denoting the cotangent space at $x$. Lie derivatives along a vector field $v$ are denoted $\mathscr{L}_{v}$.
@@ II.1 One-Particle Distribution Function and Its Invariance
In a spacetime $(M,g)$, the dynamical state of a particle with mass $m$ is specified by a position $x\in M$ and a physical momentum $p_{a}\in T^{*}_{x}M$. The future mass shell is
$\Gamma^{+}_{m}\coloneqq\left\{(x,p_{a})\in T^{*}M\mid x\in M,\ p_{a}\in F^{+}{x}(m)\right\},$ (1)
@@
$f\colon\Gamma^{+}\to\mathbb{R},$ (3)
where $\Gamma^{+}\equiv\displaystyle\mathop{\cup}_{m>0}\Gamma^{+}_{m}$. The particle current density $J_{\mu}$ and energy-momentum tensor $T_{\mu\nu}$ are
$J_{\mu}|_{x}$ $\coloneqq\int_{F^{+}_{x}}p_{\mu}f\mathrm{dvol}_{x}(p),$ $T_{\mu\nu}|_{x}$ $\coloneqq\int_{F^{+}_{x}}p_{\mu}p_{\nu}f\mathrm{dvol}_{x}(p),$ (4)
where $\mathrm{dvol}_{x}(p)$ represents the volume element on $F^{+}_{x}\equiv\displaystyle\mathop{\cup}_{m>0}F^{+}_{x}(m)$.
For free particles (influenced solely by the spacetime background) with mass $m$ and charge $q$ in Kerr-Newman (KN) spacetime with an electromagnetic field $F_{ab}=(\mathrm{d}A)_{ab}$, the equations of motion for a trajectory $C(\tau)$ are
@@
$\hat{C}(\tau)\coloneqq\left\{(x^{\mu}(\tau),p_{\nu}(\tau))\mid p_{\mu}(\tau)\equiv mg_{\mu\nu}|_{x(\tau)}T^{\nu}(\tau)\right\},$ (6)
@@
Next, we consider Killing vector fields on $M$. Let $\phi\colon M\to M$ be an isometry generated by a Killing vector field $K$. The original trajectory $C(\tau)$ naturally induces a transformed trajectory $\tilde{C}(\tau)\equiv\phi\circ C(\tau)$, with tangent vector $\tilde{T}^{c}|_{\phi(x)}=\phi_{*}[T^{c}|_{x}]$, which preserves the dynamical equations due to the invariance properties:
@@
Beyond these symmetry considerations, we impose no additional constraints on the distribution function $f\colon\Gamma^{+}\to\mathbb{R}$.
@@ II.2 Distribution Function in the Phase Space
where $\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\varDelta=r^{2}-2\mathcal{M}r+a^{2}+Q^{2},\varSigma^{2}=(r^{2}+a^{2})^{2}-\varDelta a^{2}\sin^{2}\theta$ and $a=\frac{J}{M}$. The electromagnetic vector potential is given by
@@
The Killing vectors $\xi^{a}=\left(\frac{\partial\ }{\partial t}\right)^{a},\psi^{a}=\left(\frac{\partial\ }{\partial\varphi}\right)^{a}$ and Killing tensorsCarter(1973) ; Wald(1984) $g_{ab},K_{ab}\equiv 2\rho^{2}l_{(a}k_{b)}+r^{2}g_{ab}$ (where $l^{a}\equiv\frac{1}{\ \Delta\ }\left[(r^{2}+a^{2})\xi^{a}+a\psi^{a}+\Delta\left(\frac{\partial\ }{\partial r}\right)^{a}\right],k^{a}\equiv\frac{1}{2\rho^{2}}\left[(r^{2}+a^{2})\xi^{a}+a\psi^{a}\right.$ $\left.-\Delta\left(\frac{\partial\ }{\partial r}\right)^{a}\right]$) provide the constants of motion
$E$ $\coloneqq-\pi_{a}\xi^{a},$ $L_{z}$ $\coloneqq\pi_{a}\psi^{a},$ $H$ $\coloneqq\frac{1}{2}g_{ab}p^{a}p^{b},$ $L^{2}$ $\coloneqq K_{ab}p^{a}p^{b},$ (12)
@@
Using these constants of motion as the canonical momentum coordinates,
@@
we define the conjugate configuration variables via
$Q^{\mu}\coloneqq\frac{\partial S}{\partial P_{\mu}},$ (14)
@@
where $\pi(K)\equiv\left[\pi_{\mu}(\mathrm{d}x^{\mu})_{A}\right]\left[K^{\mu}\left(\frac{\partial\ }{\partial x}^{\mu}\right)^{A}\right]$ and $K^{a}\equiv K^{\mu}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{a}$ is a Killing vector on $M$. This yields the explicit form of the Vlasov equation in Eq. (7):
@@ III Particle Trajectory Analysis
The equations of motion, Eq. (12), for charged particles in KN spacetime are
@@
where $\kappa\equiv qQ$. To facilitate subsequent discussions, we define several dimensionless parameters:
@@
The angular motion Eq. (20b) gives the following parametrized system
@@
where $\bar{L}\equiv\sqrt{L^{2}-a^{2}\cos^{2}\theta}$. The radial motion Eq. (20a) can be reformulated as
@@
thus $E\geq 1$ is necessary.
At a given spatial position $(r,\theta)$, the angular motion Eq. (20b) requires $\bar{L}\geq 0$, while the radial motion Eq. (24) demands $R(r)\geq 0$. These constraints define the parameter ranges $D_{\mathrm{abs}}$ for particles absorbed by the black hole and $D_{\mathrm{scat}}$ for scattered particles:
$D_{\mathrm{abs}}$ $\equiv D_{\sigma}\times D_{E\,\mathrm{abs}}\times D_{\bar{L}\,\mathrm{abs}}=[0,2\pi]\times[1,+\infty)\times(0,\bar{L}_{c}),$ (26a)
$D_{\mathrm{scat}}$ $\equiv D_{\sigma}\times D_{E\,\mathrm{scat}}\times D_{\bar{L}\,\mathrm{scat}}=[0,2\pi]\times[E_{c},+\infty)\times(\bar{L}_{c},\bar{L}_{\mathrm{max}}),$ (26b)
where the parameter $\bar{L}$ (rather than $L$) is adopted due to its convenience for studying specific spatial positions. The upper bound $\bar{L}_{\mathrm{max}}=\bar{L}_{\mathrm{max}}(r,\theta,\sigma,E)$ is determined by the $R(r,\theta,\sigma,E,\bar{L})=0$,
@@
The critical parameters $E_{c}-r$ and $\bar{L}_{c}-E$ can be determined by solving the periastron condition ($\dot{r}=0$): $R=0$ and $R^{\prime}\equiv\frac{\partial R}{\partial r}=0$. Simplifying under the condition $02rE-\kappa>$ 0 yields:
@@
Decoupling $E$ and $\bar{L}$ (involving roots of quartic equations) gives formal expressions $E=f_{E}(a,Q,r,\theta,\sigma,q)$, and $\bar{L}=f_{\bar{L}}(a,Q,r,\theta,\sigma,q)$. Numerical analysis reveals two critical radii for $r>r_{H}=1+\sqrt{1-(a^{2}+Q^{2})}$: a singularity $r_{\mathrm{ph}}=r_{\mathrm{ph}}(a,Q,\theta,\sigma)$ where $f_{E}\to+\infty$ and a root $r_{\mathrm{mb}}=r_{\mathrm{mb}}(a,Q,\theta,\sigma,q)$ where $f_{E}=1$. The monotonicity of $f_{E}-r$ over $(r_{\mathrm{ph}},r_{\mathrm{mb}})$ allows the formal derivation of $\bar{L}=g_{\bar{L}}(a,Q,\theta,E,\sigma,q)$
$E_{c}$ =$=\begin{cases}+\infty&\text{for }r\leq r_{\mathrm{ph}},\\ f_{E}(a,Q,r,\theta,\sigma,q),&\text{for }r_{\mathrm{ph}}<r<r_{\mathrm{mb}},\\ 1,&\text{for }r\geq r_{\mathrm{mb}}.\end{cases},$ $\bar{L}_{c}$ =$=\begin{cases}g_{\bar{L}}(a,Q,\theta,E,\sigma,q)&\text{for }E>1;\\ \bar{L}_{\mathrm{mb}}&\text{for }E=1.\end{cases}$ (29)
where $\bar{L}_{\mathrm{mb}}\equiv f_{\bar{L}}|_{r=r_{\mathrm{mb}}}$. Numerical examples are shown in Fig. 1, while analytical derivations of $E_{c}$ and $\bar{L}_{c}$ are detailed in Appendix B.
@@
Figure 1: For a KN black hole ($M=1$, $a=0.1$, $Q=0.1$), critical parameters (a) $E_{c}$ as a function of periastron radius $r$ and (b) $L_{c}$ as a function of $E$ on the equatorial plane ($\theta=\pi/2$) for various $\kappa$, with $\sigma=\pi/2$.
Fig. 1a shows $E_{c}$ as a function of periastron radius $r$. For fixed $\kappa$, $E_{c}(r)$ has a vertical asymptote at $r=r_{\mathrm{ph}}$ and $E_{c}(r_{\mathrm{mb}})=1$. When $r\in(r_{\mathrm{ph}},r_{\mathrm{mb}})$, $E_{c}$ increases as $r$ decreases. No scattering occurs for $r<r_{\mathrm{ph}}$, while any energy $E_{c}\geq 1$ permits scattering when $r>r_{\mathrm{mb}}$. Fig. 1b shows that $L_{c}$ grows approximately linearly with $E$ for fixed $\kappa$, with curves shifting downward as $\kappa$ increases.